OpenStudy (anonymous):

Write an equation of a line with a slope of 0 and passing through the point (5,4).

OpenStudy (anonymous):

\[y-4=0(x-5),y-4=0,y=4\]

OpenStudy (anonymous):

it's basically just a straight line 4 spaces above the x-axis, since it has no slope and has to pass 4....

OpenStudy (anonymous):

thank you!